# IPO: Iterative Preference Optimization for Text-to-Video Generation

### v2

@@ 4.3 Iterative training pipeline
Iterative KTO
KTO [8] is an offline learning method that leverages pointwise scoring data, eliminating the need for preference-based datasets. By directly utilizing binary-labeled data, KTO effectively trains algorithms while demonstrating heightened sensitivity to negative samples. This makes it particularly adept at handling imbalanced datasets with uneven distributions of positive and negative samples. Furthermore, KTO achieves significant performance improvements even without relying on the SFT stage. In scenarios where data is imbalanced, paired sorting data is unavailable, or preference data is noisy, KTO stands out as a more effective and robust solution.
  • Video Generation. Similar to the DPO training framework, we adopt an identical video sampling strategy in this study. In alignment with the KTO [8] methodology, empirical evidence suggests that sampling multiple videos for a given prompt yields superior results compared to single-video sampling. Therefore, we maintain consistency with the DPO sampling protocol throughout our implementation.
  • Pointwise scoring. The Critic model is employed to evaluate and assign scores to all generated text-video pairs $(y,x)$. Each data sample is categorized into three distinct quality levels: ”Good”, ”Normal”, and ”Bad”. Based on this classification, we designate samples labeled as ”Good” and ”Normal” as positive instances, while those marked as ”Bad” are treated as negative instances for subsequent analysis and model training.
  • Preference learning. We use the Eq. 4 for training. KTO needs to estimate the average reward Q during training. When the batch size is small, the Q value is not accurately estimated and the model training will be unstable. We set the batch size to 128 and achieved good results. To stabilize the training, we also sampled data from the real data VidGen-1M [42] and added negative log-likelihood loss.
    $\mathcal{L}(\theta)$ $=L_{diffusion-kto}(\theta)$ (6)
    $\quad+\lambda\cdot\mathbb{E}_{(\textbf{{y}},x)\sim\mathcal{D}^{\mathrm{real}}}\bigl{[}-\log p_{\theta}(y|x)\bigr{]},$
Both DPO and KTO employ an iterative learning framework that cyclically executes three core steps: (1) strategic data sampling, (2) model-based preference labeling, and (3) alignment optimization. The process incorporates early stopping based on validation metrics while adaptively optimizing strategies for each component to maximize training efficiency and model performance.
@@ 5 Experiments
KTO 82.47 83.56 78.08 96.85 97.08 99.27 97.93 62.47 68.83 62.31
DPO 82.42 83.53 77.97 96.81 97.23 99.21 98.03 62.35 68.76 61.77
Table 1: Verify the universality and effectiveness of iterative preference optimization, evaluate the pairwise-based DPO and the KTO algorithm based on pointwise data.
@@
IPO-iter1 81.69 82.87 77.01 96.77 97.01 99.01 97.84 61.55 64.33 62
IPO-iter2 82.42 83.53 77.97 96.81 97.23 99.21 98.03 62.35 68.76 61.77
IPO-iter3 82.74 83.92 78.00 96.79 97.48 99.35 98.17 62.31 69.46 62.87
Table 2: Verify the effect of iterative preference optimization and multiple rounds of training, we evaluated the performance after three rounds of iterative optimization on VBench
Models
Total
Score.
Quality
Score.
Semantic
Score.
Subject
Consist.
Background
Consist.
Temporal
Flicker.
Motion
Smooth.
Aesthetic
Quality
Dynamic
Degree
Image
Quality
vila 13b 82.42 83.53 77.97 96.81 97.23 99.21 98.03 62.35 68.76 61.77
vila 40b 82.52 8363 7807 96.83 97.37 99.19 98.09 62.38 68.92 62.01
@@
Figure 2: Qualitative Comparison. We compare our (iterative preference optimization) performance with CogVideoX-2B.
This section presents the experimental setup, results, and ablation studies to demonstrate the effectiveness of the proposed online reinforcement learning framework for video generation optimization using human feedback during post-training.
5.1 Experimental Setup
Our experiments are conducted on a Critic model constructed based on the diffusion model, which evaluates video quality through semantic and temporal consistency. The Critic model incorporates both metrics derived from human feedback and pre-defined objective criteria to ensure comprehensive evaluation.
The training process leverages both DPO and KTO methods within the proposed online reinforcement learning framework. The DPO method aligns the generated videos with human feedback by minimizing divergence between the learned policy and a reference policy, while KTO directly maximizes the utility of the generated video instead of maximizing the preference log-likelihood. Both methods were implemented with a batch size of 128, learning rate of 2e-5, and trained for up to three rounds of iterative optimization. Each round involves a combination of human feedback-guided rewards and the original diffusion model loss to balance consistency and realism.
5.2 Results
5.2.1 Quantitative Results
Table 4: Comparison of the proposed method with state-of-the-art (SOTA) approaches. Metrics are evaluated using VBench.
Method Quality Score Semantic Score Total Score
Baseline $2B$ 82.18 75.83 80.91
Baseline $5B$ 82.75 77.04 81.61
Optimized $2B$ (Ours) 83.92 78.00 82.74
Models
Subject
Consist.
Background
Consist.
Temporal
Flicker.
Motion
Smooth.
Aesthetic
Quality
Dynamic
Degree
Image
Quality
CogVideoX-2B 96.78 96.63 98.89 97.73 60.82 59.86 61.68
w/o sft loss 96.56 96.83 98.73 97.81 61.25 60.94 61.33
w/ sft loss 96.77 97.01 99.01 97.84 61.55 64.33 62
Table 5: Performance comparison with and without real data training.To verify which model training is more stable, we compared experiments with and without the inclusion of real data. Metrics are evaluated using VBench.
Quantitative results are evaluated using the VBench metric, which comprehensively measures video generation quality in terms of temporal coherence, semantic alignment, and overall plausibility. Our experiments demonstrate that the optimized $2B$ model achieves significant improvements across all evaluated metrics compared to the baseline $5B$ model. Specifically, the optimized model exhibits substantial gains in temporal coherence, effectively enhancing the smoothness of transitions between frames and minimizing abrupt changes. Moreover, the model shows marked improvements in semantic alignment, demonstrating a stronger ability to adhere to the intended meaning of input prompts and better capturing their underlying context. These advancements highlight the efficacy of the proposed post-training strategy in addressing key limitations of the baseline model. The overall VBench score further confirms that the optimized $2B$ model not only achieves superior video generation quality but also maintains computational efficiency, underscoring the practical advantages of our approach in optimizing smaller-scale models to outperform larger counterparts.
5.2.2 Qualitative Results
To further analyze the benefits of our approach, we provide qualitative comparisons between the baseline and optimized models. Visual results demonstrate that the optimized $2B$ model produces videos with higher semantic fidelity, accurately capturing the intent of input prompts and reflecting their underlying context. Additionally, the optimized model ensures greater temporal consistency, effectively minimizing artifacts and sudden transitions between frames, which are often observed in the baseline outputs. Furthermore, the generated videos exhibit enhanced realism, with reduced incoherent structures and improved plausibility, making the outputs more visually appealing and aligned with human perception. These qualitative improvements highlight the effectiveness of our optimization process in addressing key challenges related to semantic and temporal consistency, significantly elevating the overall quality of generated videos.
5.3 Ablation Studies
We conduct a series of ablation studies to validate the contributions of key components in our framework. Each experiment evaluates performance using the VBench metric to ensure consistency in evaluation.
Model Scale Pairwise Pointwise
Accuracy Accuracy
13B 83.72 95.33
40B 86.14 97.57
5.3.1 Impact of Critic Model Scale on Accuracy
To evaluate the impact of model scale on the performance of the critic model, we conducted experiments with two different scales of ViLA-based critic models: 13B and 40B parameters. The critic models were fine-tuned on a combination of pairwise and pointwise annotated datasets, designed to align with human preferences effectively. Table 6 presents the accuracy of these models on validation datasets for both pairwise ranking and pointwise scoring tasks. The results indicate that increasing the model scale from 13B to 40B leads to significant improvements in both pairwise and pointwise accuracy. Specifically, the 40B model achieves 86.14% accuracy in pairwise ranking and 97.57% in pointwise scoring, outperforming the 13B model by 2.42% and 2.24%, respectively. This demonstrates that larger-scale critic models are better at capturing and generalizing human preferences, likely due to their increased capacity to represent complex relationships in multimodal data. The ablation study highlights the importance of model scale in improving the critic model’s performance. While smaller models can achieve reasonable results, larger models provide a notable boost in accuracy, which is crucial for downstream reinforcement learning tasks. These findings suggest that scaling up the critic model can be a straightforward yet effective way to enhance its alignment capabilities and improve the overall quality of video generation.
5.3.2 Incorporating Real Video Data
As shown in Table 5, the introduction of real video data (indicated by the inclusion of sft loss) leads to measurable improvements across multiple VBench metrics, including temporal flicker, motion smoothness, and aesthetic quality. Notably, the optimized model achieves higher scores in semantic consistency and dynamic degree, indicating a significant enhancement in the realism and coherence of generated videos. These results underscore the effectiveness of integrating real-world data with reinforcement learning to refine video generation outputs, enabling the model to produce videos with superior visual and temporal fidelity.
5.3.3 Comparison of different RL Methods
To evaluate the performance of the proposed online reinforcement learning framework, we conducted experiments using two different feedback alignment methods: Knowledge Transfer Optimization (KTO) and Direct Policy Optimization (DPO). The results, presented in Table 1, demonstrate that our framework consistently improves video generation quality across multiple metrics, regardless of the specific reinforcement learning method employed.
Between the two methods, KTO achieves consistently better performance than DPO across most metrics. Notably, KTO yields a higher Dynamic Degree score (68.83 vs. 68.76) and improved Motion Smoothness (97.93 vs. 98.03), suggesting that KTO is more effective at producing realistic and temporally coherent videos. Additionally, KTO demonstrates greater stability during training, as evidenced by its higher overall scores across Subject Consistency and Background Consistency, which measure adherence to semantic and contextual details. These findings highlight the advantages of the KTO method in leveraging multi-step optimization, enabling the generation of more plausible and visually consistent video outputs compared to DPO.
In summary, the experimental results confirm that the proposed online reinforcement learning framework is effective under different optimization methods. Among these, KTO exhibits superior performance and stability, making it a preferred choice for improving video generation quality through human feedback alignment.
5.3.4 Effect of Multi-Round Training
The impact of multi-round reinforcement learning is assessed by training models with 1, 2, and 3 rounds of iterative optimization, where a single round represents the traditional offline post-training approach. Results in Table 2 show that performance improves with each additional round, as multi-round training better aligns the model with human feedback and refines generation quality. These findings confirm that iterative reinforcement learning effectively enhances both temporal and semantic consistency.
6 Conclusion
This paper presents Iterative Preference Optimization (IPO), a novel post-training framework for text-to-video (T2V) generation. IPO iteratively enhances models using preference datasets, criticism models, and iterative learning, improving subject consistency, motion smoothness, and aesthetic quality. As the first to apply iterative preference optimization to T2V, IPO leverages human feedback to refine video quality, making it more aligned with user expectations. Experiments show that IPO enables a 2B-parameter model to outperform a 5B-parameter model on VBench.

### v3

@@ 3.3.3 Pairwise Optimization
where $\mathcal{D}^{\mathrm{sample}}$ represents the set of preference pairs derived from the critic model’s evaluation, and $\mathcal{D}^{\mathrm{real}}$ denotes the real video data used for regularization.
3.3.4 Iterative Optimization
Iterative optimization is advantageous for complex tasks like preference alignment in reinforcement learning. Rather than approximating the optimal policy $\pi_{\theta}$ in a single step, which is computationally challenging, it decomposes the problem into manageable stages. This stepwise refinement enables incremental improvements, progressively aligning the model with human preferences in a more efficient and controlled manner. A key benefit of iterative optimization is that it enables more stable training. Rather than attempting to directly approximate the reference distribution $\pi_{\text{ref}}\exp\left(\frac{1}{\beta}r(x_{0},c)\right)$ in a single step, the process proceeds in multiple stages. In each iteration, we approximate the distribution using the following formula:
$\pi_{\text{ref}}^{(k)}=\pi_{\text{ref}}^{(k-1)}\exp\Big{(}\frac{r^{k}}{\beta_{k}}\Big{)},$ (10)
where, $r^{k}$ denotes the reward function at the $k$-th iteration, and $\beta_{k}$ is a hyperparameter controlling the optimization rate at each stage. We initialize $r^{0}=0$ to start with a neutral reward function. As iterations progress, the reward function is refined, guiding the model toward an optimal policy.
Iterative refinement improves training control, enabling the model to gradually align with human feedback while avoiding early overfitting or divergence. By utilizing intermediate distributions at each stage, the method accommodates more complex feedback signals, leading to a more robust model that generalizes well across diverse real-world scenarios. Additionally, iterative optimization mitigates issues like catastrophic forgetting and instability, which can occur when attempting to approximate the optimal policy in a single step. Each iteration builds on the previous one, ensuring stable and reliable convergence toward the target policy. In this way, iterative preference optimization (IPO) enhances preference-based optimization by providing better training control, faster convergence, and improved generalization, making it a valuable tool for optimizing policies based on human feedback.
Method Quality Semantic Total
Score Score Score
CogVideoX-2B 82.18 75.83 80.91
CogVideoX-5B 82.75 77.04 81.61
CogVideoX-2B-IPO-3 83.92 78.00 82.74
Table 1: Comparison of IPO with baselines CogVideoX models.
Models Subject Motion Aesthetic
Consist. Smooth Quality
CogVideoX-2B 96.78 97.73 60.82
w/o real data sft 96.56 97.81 61.25
w/ real data sft 96.77 97.84 61.55
Table 2: Ablation studies on effects of real data for finetuning IPO.
@@ 4 Experiments
CogVideoX-2B-IPO1 81.69 82.87 77.01 96.77 97.01 99.01 97.84 61.55 64.33 62
CogVideoX-2B-IPO2 82.42 83.53 77.97 96.81 97.23 99.21 98.03 62.35 68.76 61.77
CogVideoX-2B-IPO3 82.74 83.92 78.00 96.79 97.48 99.35 98.17 62.31 69.46 62.87
Table 3: Ablation studies on the iteration number adopted in IPO. The evaluation is performed on VBench.
Figure 4: Qualitative Comparison of our proposed IPO model and the baseline CogVideoX-2B model. We can see IPO improves the baseline in the aspects of prompt following (the 1st line), aesthetic quality (the 2nd line), subject consistency (the 3rd line) and motion smoothness (the 4th line). This demonstrates the efffectiveness of our IPO for enhancing the generation capability of video found models. Additional visualization results are available in the supplementary materials.
4.1 Experimental Setup
4.1.1 Critic Model Training
To unify paired data sorting and individual video scoring, we fine-tune the pre-trained VILA [19], leveraging its strengths in video comprehension and instruction following. Key modifications include integrating reasoning chains before scoring or ranking and enhancing decision-making with explicit thought processes. This improves result accuracy and the quality of natural language feedback.
[…]
4.1.2 T2V Alignment Training
[…]
4.2 Results
4.2.1 Quantitative Results
Quantitative results are evaluated using the VBench metric, which measures video generation quality in terms of temporal coherence, semantic alignment, and overall plausibility. As shown in Tab. 1, our experiments reveal that the optimized $2B$ model outperforms the baseline $5B$ model across all metrics. The optimized model shows significant improvements in temporal coherence, enhancing frame transitions and minimizing abrupt changes. It also demonstrates stronger semantic alignment, better capturing the intended meaning and context of input prompts. These results highlight the effectiveness of our post-training strategy in addressing key limitations of the baseline model. The overall VBench score further confirms that the optimized $2B$ model achieves superior video quality while maintaining computational efficiency, making it a practical solution for optimizing smaller-scale models.
4.2.2 Qualitative Results
To further assess the benefits of our approach, we provide qualitative comparisons between the baseline and optimized models. Visual results, shown in Fig. 4, reveal that the optimized $2B$ model generates videos with better semantic fidelity, accurately capturing the intent of input prompts and their context. It also ensures improved temporal consistency by minimizing artifacts and sudden transitions between frames, common in the baseline outputs. Additionally, the optimized videos exhibit enhanced realism, with fewer incoherent structures and greater plausibility, making them more visually appealing and aligned with human perception. This further demonstrate the effectiveness of IPO.
4.3 Ablation Studies
4.3.1 Incorporating Real Video Data
As shown in Tab. 2, the introduction of real video data (indicated by the inclusion of sft loss) leads to measurable improvements across multiple VBench metrics, including temporal flicker, motion smoothness, and aesthetic quality. Notably, the optimized model achieves higher scores in semantic consistency and dynamic degree, indicating a significant enhancement in the realism and coherence of generated videos. These results underscore the effectiveness of integrating real-world data with reinforcement learning to refine video generation outputs, enabling the model to produce videos with superior visual and temporal fidelity.
Method Quality Semantic Total
Score Score Score
CogVideoX-2B 80.91 82.18 75.83
CogVideoX-2B-KTO1 82.47 83.56 78.08
CogVideoX-2B-DPO1 82.42 83.53 77.97
Table 4: Ablation studies on impacts of different preference optimization strategies for IPO: DPO and KTO.
4.3.2 Comparison of different RL Methods
To evaluate the performance of the proposed online reinforcement learning framework, we conducted experiments using two different feedback alignment methods: Knowledge Transfer Optimization (KTO) and Direct Policy Optimization (DPO). The results, presented in Tab. 4, demonstrate that our framework consistently improves video generation quality across multiple metrics, regardless of the specific reinforcement learning method employed. All evaluation metrics across different dimensions are provided in the supplementary materials, further supporting this comprehensive improvement.
Between the two methods, KTO achieves consistently better performance than DPO across most metrics. Notably, KTO yields a higher Dynamic Degree score (68.83 vs. 68.76) and improved Motion Smoothness (97.93 vs. 98.03), suggesting that KTO is more effective at producing realistic and temporally coherent videos. Additionally, KTO demonstrates greater stability during training, as evidenced by its higher overall scores across Subject Consistency and Background Consistency, which measure adherence to semantic and contextual details. These findings highlight the advantages of the KTO method in leveraging multi-step optimization, enabling the generation of more plausible and visually consistent video outputs compared to DPO. This confirms our IPO is flexible to adapt to different optimization methods. In addition, KTO exhibits superior performance and stability, making it a preferred choice.
4.3.3 Effect of Multi-Round Training
The impact of multi-round reinforcement learning is assessed by training models with 1, 2, and 3 rounds of iterative optimization, where a single round represents the traditional offline post-training approach. Results in Tab. 3 show that performance improves with each additional round, as multi-round training better aligns the model with human feedback and refines generation quality. These findings confirm that iterative reinforcement learning effectively enhances both temporal and semantic consistency.
Figure 5: Human evaluation of IPO and CogVideoX-2B.
4.4 Human Evaluation
Fig. 5 presents the results of a human evaluation comparing the IPO model with the baseline CogVideo-2B The result shows that our IPO outperforms the baseline in all dimensions. In the ”Consistency” dimension, the two models performed equally in around 40.5 percent of the cases, indicating that the baseline excels in maintaining text consistency. However, in the ”Motion” and ”Faithfulness” dimensions, the baseline’s performance is average, while the IPO model demonstrates a significant improvement, particularly in generating more natural motion and better alignment with the given text. All evaluation metrics across different dimensions are provided in the supp. materials
5 Conclusion
In this paper, we present a novel Iterative Preference Optimization framework for improvement video foundation models from the pespective of post training. In particular, IPO collects a human preference dataset and trains a critic model for automatically labeling the generated videos from base model. Given this, IPO efficiently optimizes the base model in an iterative manner. By this way, IPO can effectively align the base model with human preference, leading to enhanced video results in motion smoothness, subject consistency and aesthetic quality, etc. In addition, we adapt IPO with different optimization strategies: DPO and KTO, both achieve impressive improvements. This further verifies the flexibility of IPO. Thorough experiments on VBench demonstrates the efficacy of IPO.
Supplementary Material
Figure 6: Our user interface presents one sample at a time to annotators, featuring four distinct icons for different functions.
Appendix A Human Annotation
Prior to the final manual annotation, we undertook several steps to ensure the quantity, quality, and efficiency of the annotation process. In building the dataset, we focused on the diversity and complexity of prompts, creating annotated data across multiple dimensions and categories. We meticulously selected candidates for the annotation task and provided them with in-depth training sessions. We then designed a user-friendly interface and a comprehensive annotation program to streamline the process. Additionally, we developed a detailed annotation guide that covers all aspects of the task and highlights essential precautions. To further ensure consistency and accuracy among annotators, we conducted multiple rounds of annotation and implemented random sampling quality checks.
A.1 Annotator selection
The effectiveness of annotated data hinges significantly on the capabilities of the individuals responsible for the annotation. As such, at the onset of the annotation initiative, we placed a strong emphasis on meticulously selecting and training a team of annotators to ensure their competence and objectivity. This process entailed administering a comprehensive evaluation to potential candidates, which was designed to gauge their proficiency in ten critical domains: domain expertise, tolerance for visually disturbing content, attention to detail, communication abilities, reliability, cultural and linguistic competence, technical skills, ethical awareness, aesthetic discernment, and motivation.
Given that the models under assessment could generate videos containing potentially distressing or inappropriate material, candidates were informed of this possibility in advance. Their participation in the evaluation was predicated on their acknowledgment of this condition, and they were assured of their right to withdraw at any time. Based on the results of the evaluation and the candidates’ professional backgrounds, we sought to form a team that was diverse and well-rounded in terms of expertise and capabilities across the ten assessed areas. Ultimately, we selected a group of 10 annotators—comprising five men and five women, all of whom possessed bachelor’s degrees. We conducted follow-up interviews with the selected annotators to reaffirm their suitability for the annotation task.
A.2 Annotation Guideline
When it comes to assessing human preferences in videos, the evaluation process is structured around three fundamental aspects: semantic consistency, motion smoothness, and video faithfulness. Each of these dimensions is meticulously examined independently to provide a comprehensive understanding of the video’s overall quality. Semantic consistency ensures that the video’s content aligns logically and contextually, while motion smoothness evaluates the smoothness and naturalness of movements within the video. Video faithfulness, on the other hand, assesses how well the video maintains a realistic and high-quality appearance. By breaking down the evaluation into these distinct yet interconnected aspects, we can achieve a more nuanced and accurate assessment of human preferences in video content.
  • Semantic Consistency The consistency between the text and the video refers to whether elements such as the topic category (e.g., people or animals), quantity, color, scene description, and style mentioned in the text align with what is displayed in the video.
  • Motion Smoothness The motion smoothness refers to whether the movements depicted are smooth, coherent, and logically consistent, particularly in terms of their adherence to physical laws and real-world dynamics. This includes evaluating if the actions unfold in a seamless and believable manner, without abrupt or unrealistic transitions, and whether they align with the expected behavior of objects or individuals in the given context.
  • Video Faithfulness Faithfulness refers to the accuracy and realism of the visual content in the video. It examines whether depictions of people, animals, or objects closely match their real-world counterparts, without unrealistic distortions like severed limbs, deformed features, or other anomalies. High fidelity ensures the visual elements are believable and consistent with reality, enhancing the video’s authenticity..
Pointwise Annotation
For each dimension, annotators assign one of three ratings:
  • Good: No observable flaws; fully satisfies the criterion.
  • Normal: Minor imperfections that do not severely impact perception.
  • Bad: Significant violations that degrade quality or realism.
Annotations are performed independently across dimensions to avoid bias.
Pairwise Annotation
Given two videos $\mathcal{V}_{A}$ and $\mathcal{V}_{B}$, annotators compare them under each dimension:
  • For semantic consistency, determine which video better aligns with the intended context.
  • For motion smoothness, identify the video with more natural and continuous motion.
  • For video faithfulness, select the video exhibiting higher realism and fewer artifacts.
A ranked preference ($\mathcal{V}_{A}\succ\mathcal{V}_{B}$ or $\mathcal{V}_{B}\succ\mathcal{V}_{A}$) is recorded separately for each dimension. Ties are permitted only if differences are indistinguishable.
A.3 Annotation training and interface
We convened a comprehensive training session centered on our detailed user guidelines to ensure that the annotation team fully comprehends our goals and standards. During the training, we covered the purpose, precautions, standards, workload, and remuneration related to the annotation process. Additionally, we formally communicated to the annotators that the human annotation is strictly for research purposes and that the annotated data might be made publicly available in the future. We reached a mutual agreement with the annotators regarding the standards, workload, remuneration, and intended use of the annotated data. The criteria for assessing video faithfulness, text-video alignment, and motion smoothness are universal, which means that individual standards should not vary significantly. Consequently, we annotated several sample videos using our carefully designed annotation platform to ensure consistency among annotators. An overview of the developed annotation platform is shown in Fig. 6. Through this training, we also equipped the annotators with the necessary knowledge to conduct impartial and detailed human evaluations of video faithfulness, text-video alignment, and motion smoothness.
Model Subject Consistency Background Consistency Temporal Flickering Motion Smoothness Dynamic Degree Aesthetic Quality Imaging Quality Object Class
CogVideoX-2B 96.78 96.63 98.89 97.73 59.86 60.82 61.68 83.37
CogVideoX-2B-KTO1 96.77 97.01 99.01 97.84 64.33 61.55 62 85.41
CogVideoX-2B-KTO2 96.81 97.23 99.21 98.03 68.76 62.35 61.77 85.63
CogVideoX-2B-KTO3 96.79 97.48 99.35 98.17 69.46 62.28 62.87 85.77
Model Multiple Objects Human Action Color Spatial Relationship Scene Appearance Style Temporal style Overall Consistency
CogVideoX-2B 62.63 98.00 79.41 69.90 51.14 24.8 24.36 26.66
CogVideoX-2B-KTO1 63.25 98.54 81.67 67.83 54.15 25.13 24.86 27.03
CogVideoX-2B-KTO2 63.44 98.97 83.24 70.16 54.35 25.47 26.44 27.37
CogVideoX-2B-KTO3 63.51 99.03 82.33 70.23 54.41 25.48 25.39 27.66
Table 5: Quantitative results on video assessment metrics. Display of comparison results between all indicators and baseline on vbench through multiple iterations.
Appendix B Additional Quantitative Analysis
B.1 All evaluation dimension results
We present the evaluation metrics of VBench benchmark in 16 different dimensions in the Tab. 5. The results indicate that the performance shows a continuous upward trend in multiple training iterations, indicating that the training is dynamically stable. Although small fluctuations are observed, they can be ignored and will not affect the overall positive progress of the indicator.
Model Scale Pairwise Pointwise
Accuracy Accuracy
IPO-Critic-13B 78.41 82.95
IPO-Critic-40B 81.33 85.57
B.2 Impact of Critic Model Scale on Accuracy
To evaluate the impact of model scale on critic performance, we experimented with two ViLA-based critic models of 13B and 40B parameters. Fine-tuned on paired and pointwise annotated datasets, these models were assessed on validation tasks, as shown in Tab. 6.
Results show that increasing the model size from 13B to 40B significantly improves accuracy, with the 40B model achieving 80.73% in paired ranking and 86.57% in pointwise scoring—2.92% and 2.62% higher than the 13B model. This suggests that larger critic models better capture human preferences by representing complex multimodal relationships.
Our ablation study further highlights the importance of model scale. While smaller models perform reasonably well, larger models significantly boost accuracy, benefiting downstream reinforcement learning tasks. These findings suggest that scaling the critic model is a simple yet effective way to enhance preference alignment and improve video generation quality.
@@ B.2 Impact of Critic Model Scale on Accuracy
w/IPO-Critic-13B 82.42 83.53 77.97 96.81 97.23 99.21 98.03 62.35 68.76 61.77
w/IPO-Critic-40B 82.52 8363 78.07 96.83 97.37 99.19 98.09 62.38 68.92 62.01
@@
B.3 Impact of Critic Model Size on RL
We conducted a comparative analysis of critic models of different sizes in the iterative preference optimization training process. As shown in Tab. 7, larger critic models provide more precise reward estimations, leading to improved alignment with human preferences. This enhanced reward accuracy contributes to better optimization during preference alignment training, ultimately resulting in superior model performance. Our findings highlight the importance of scaling the critic model to achieve more reliable preference optimization and reinforce the benefits of leveraging larger critic architectures for improved alignment outcomes.
Appendix C Additional Qualitative Comparison
We provide a more detailed qualitative comparison in Fig. 7, which demonstrates the superiority of our model over the baseline. This visualization highlights key improvements in fidelity and overall quality.By comparing outputs side by side, we illustrate the advantages of our model in handling complex scenarios, reinforcing its effectiveness in real-world applications.
Appendix D Societal Impacts
The emergence of advanced text-to-video generation pipelines and optimized training strategies marks a major breakthrough in AI-driven content creation. These innovations significantly lower barriers to video production, making it more efficient, cost-effective, and widely accessible. Industries such as education, entertainment, marketing, and accessibility technology stand to benefit immensely, as creators can generate high-quality videos with minimal effort and resources.
However, alongside these advancements come ethical concerns, particularly regarding potential misuse. One of the most pressing risks is the creation of highly realistic yet misleading content, such as deepfakes, which could be exploited for misinformation, reputational harm, or social manipulation. Additionally, automated video curation may unintentionally reinforce biases present in training datasets, leading to the perpetuation of stereotypes or discriminatory narratives.
On a larger scale, the widespread adoption of AI-driven video generation could disrupt traditional media industries, impacting professions like video editing, animation, and scriptwriting. The proliferation of hyper-realistic AI-generated content also raises concerns about media authenticity, making it increasingly difficult for audiences to distinguish between real and synthetic footage. Furthermore, if access to these technologies remains limited to well-funded organizations, existing digital divides could widen, leaving smaller creators at a disadvantage.
To mitigate these risks, proactive measures must be taken. Implementing transparency features—such as metadata labels to distinguish AI-generated content—can enhance accountability. Ensuring diverse and ethically curated datasets is crucial to reducing bias and promoting fairness in outputs. Policymakers, industry leaders, and researchers must work together to develop regulatory frameworks that balance innovation with ethical responsibility. Additionally, providing educational resources and affordable access to AI tools can help democratize their benefits, preventing technological disparities.
The rapid evolution of generative AI necessitates ongoing ethical oversight. Continuous evaluation of its societal impact, coupled with adaptive safeguards, is essential to minimizing harm while maximizing positive outcomes. Addressing these challenges requires a multidisciplinary approach, fostering a responsible AI ecosystem that empowers creators without compromising ethical integrity.
Appendix E Limitations
While our approach represents a significant advancement in text-to-video generation, several challenges persist. Addressing these issues presents valuable opportunities for future research and development. Below, we outline key limitations and potential solutions:
Computational Constraints
Despite improvements in training efficiency, large-scale text-to-video models still demand substantial computational power. This limitation may hinder accessibility, particularly for individuals and organizations with limited resources. Future research should focus on developing lightweight architectures, model compression techniques, and efficient inference strategies to enhance accessibility in resource-constrained environments.
Dataset Representativeness and Cultural Adaptability
The effectiveness of our model is highly dependent on the quality and diversity of its training dataset. While significant efforts were made to curate a representative dataset, certain cultural contexts, niche topics, and underrepresented communities may not be adequately covered. Additionally, biases present in pretraining data could lead to unintended stereotypes or inappropriate outputs. Expanding dataset coverage, improving curation techniques, and integrating fairness-aware training strategies will be crucial to enhancing contextual accuracy and mitigating bias.
Multimodal Hallucination and Interpretability
Since our evaluation models are fine-tuned multimodal large language models (MLLMs), they are susceptible to hallucination—generating content that appears plausible but is factually incorrect or cannot be inferred from the input text and images. Furthermore, MLLMs inherit biases from their pretraining data, which may lead to inappropriate responses. While careful curation of supervised fine-tuning (SFT) data helps alleviate these issues, they are not fully resolved. In addition to these concerns, MLLMs also suffer from opacity, interpretability challenges, and sensitivity to input formatting, making it difficult to understand and control their decision-making processes. Enhancing model explainability and robustness remains a crucial area for future research.
Human Annotation Limitations
Human evaluation plays a critical role in assessing model performance, but it is inherently subjective and influenced by annotator biases, perspectives, and inconsistencies. Errors or noisy labels may arise during annotation, potentially affecting evaluation reliability. To address this, we conducted multiple rounds of trial annotations and random sampling quality inspections to improve inter-annotator consistency. Additionally, we designed user-friendly annotation guidelines and interfaces to streamline the process and enhance accuracy. However, variations in annotation methodologies across different platforms and annotators can still lead to discrepancies, limiting the generalizability of our results. Moreover, human annotation remains time-consuming and resource-intensive, constraining scalability. Future efforts could explore semi-automated evaluation approaches or crowdsourced methods to improve efficiency while maintaining quality.
Ethical Considerations and Regulatory Compliance
Ensuring responsible AI deployment remains an ongoing challenge, even with safeguards in place to prevent misuse. The rapid advancement of AI-generated media demands continuous revisions to ethical policies and regulatory frameworks. Proactively engaging with policymakers, industry leaders, and researchers is crucial for developing guidelines that uphold accountability, transparency, and fair usage. Embedding metadata or digital watermarks in AI-generated content can improve traceability and help differentiate synthetic media from authentic sources. Furthermore, raising public awareness and education on AI-generated content help users recognize potential risks. Developers should prioritize responsible AI innovation by ensuring algorithmic fairness while minimizing bias. Additionally, fostering cross-industry collaboration and promoting global standards will contribute to a more secure and trustworthy AI ecosystem.
Figure 7: Enhanced qualitative comparison. We conduct a qualitative comparison between CogVideo-2B and IPO-2B to evaluate their performance